Differential Equations

6. Solve the following boundary value problem:

?? = 3???, 0 < ? < 1, ? > 0; ?(0,?) = ?(1,?) = 0; ?(?, 0) = 7 sin ?? − (1/9) sin 3?x
Answer #1

09100 ounce > U(+) = Ullt): 0 Ulm.u) = fuxo Bounduzy Condition -Initial condition. where t is tine nie tishlacement, unt) isCalenco 2 to H-WH = m²-u2 = o m 2+M ucala cenntçenia Hol - ct ucll ciene + cell Ceno cella cile EMO) 0 . C=0 (2= trivial felnHr Cn Singleton 1-IN = 64) Kolt) is (+)-dku(t)= Gilti= Aenkt -ntr- .K+ IL X Uncaut) = Hn (42. holt Un cont) = Dr. Sinnan na k# 3UXX o excl UCU. D = V1 +) = 0 U(2,0) = 7 Sinha- Kliquis k=3. here l=1 +(x)=4&inux-1 sinana then. WxH) = sin ne se va = { o), 7. Sint-I sin3t ). sinni dt S 7 D. || 유기 Sint sinnt dt - 아 Sin3t.8inni ddt 야 0 1) 11 [ Sinma. sinna 구 it mtn C 름 출명을 cind